OpenStudy (anonymous):

Solve the initial value problem for x as a function of t. (t^2 -3t+2)dx/dt =1 (t>2), x(3)=0

OpenStudy (anonymous):

integrate it?

OpenStudy (anonymous):

i have no idea. i am lost with this one

OpenStudy (anonymous):

solve for dx/dt

OpenStudy (anonymous):

and integrate

OpenStudy (anonymous):

simply dvide both ides by the (t^2-3t+2), then integrate both sides

OpenStudy (anonymous):

you will use partial fractions here

OpenStudy (anonymous):

that will get you up to x family then just use the x(3) to find the exact function

OpenStudy (anonymous):

if asnaseer dosent solve it for you i will

OpenStudy (anonymous):

ok :)

OpenStudy (asnaseer):

\[(t^2-3t+2)*\frac{dx}{dt}=1\]so:\[\int dx=\int \frac{dt}{t^2-3t+2}\]

OpenStudy (anonymous):

yes very nice^

OpenStudy (asnaseer):

this will give you:\[x=\ln(\frac{2-t}{1-t})+c\]then use the initial values you were given to find c.

OpenStudy (anonymous):

ok. thank you very much asnaseer

OpenStudy (anonymous):

i think there should be another part though, check your answer book blue12

OpenStudy (asnaseer):

yw - @Outkast3r09 basically gave the same steps but just in words.

OpenStudy (anonymous):

no i dont see another part

OpenStudy (anonymous):

well, maybe im wrong

OpenStudy (anonymous):

it is page 462. problem 51.

OpenStudy (anonymous):

you'l need to use partials to solve

OpenStudy (anonymous):

i believe he's right

OpenStudy (anonymous):

dont worry about it the intal values give you c

OpenStudy (anonymous):

yeah he is right. @Outkast3r09

OpenStudy (anonymous):

hm alright@ LagrangeSon678

OpenStudy (anonymous):

i am stuck again!. i got x=ln(2-t)/(1-t)+C how do i find C now?

OpenStudy (anonymous):

i know i am dumb..

OpenStudy (anonymous):

hy Outkast3r09 can u plz explain

OpenStudy (anonymous):

you have to put in the values for x(t)=0 you're going to get imaginary numbers i believe since inside the natural log will be negatives

OpenStudy (anonymous):

now, replace t with 0

OpenStudy (anonymous):

remeber ln(1)=0

OpenStudy (anonymous):

t with 0??

OpenStudy (anonymous):

in your anti derivaitve

OpenStudy (anonymous):

yes but i think the integration is off?

OpenStudy (anonymous):

o no its good

OpenStudy (anonymous):

why are you replacing t with 0... shouldn't it be replaced with 3?

OpenStudy (anonymous):

x=ln(t-2)-ln(t-1)+ln2 is the answer in the book

OpenStudy (anonymous):

o wait nvm i thought it was lnor ln

OpenStudy (anonymous):

ln(1/2)+c=0 c=-ln(1/2) c=-ln(1)-(-ln(2)= 0+ln(2)

OpenStudy (anonymous):

ln(t-2)-ln(t-1)=ln(t-2)/ln(t-1)

OpenStudy (anonymous):

so all together you get \[\frac{\ln(t-2)}{\ln(t-1)}+\ln(2)\]

OpenStudy (anonymous):

hm i get it thanku OutKast3r09 and LagrangeSon678

OpenStudy (anonymous):

anybody know about parametric equation

OpenStudy (anonymous):

that would hav ebeen insane if it was \[\frac{\ln(t-2)}{ln(t-1)}\] I was sitting here going .... hmm how would this be possible without imaginary... and i know smoewhat lagrange

OpenStudy (anonymous):

i know a lil about parametric equation

OpenStudy (anonymous):

can i get some help

OpenStudy (anonymous):

with?

OpenStudy (anonymous):

Find the parametric equations and a parameter interval for the motion of a particle that starts out at (a,0) and traces the x^2+y^2=a^2

OpenStudy (anonymous):

well the interval would be 0,infty due to squared i believe

OpenStudy (anonymous):

a^2 is a constant i belive lagrange?

OpenStudy (anonymous):

This is an equation of a circle with radius a.